# Align Attention Heads Before Merging Them: An Effective Way for Converting MHA to GQA

### v1

@@ 4.4 Analysis of the results
The experimental results indicate that the grouping of attention heads does have an impact on the performance of pruned models, and that grouping attention heads based on value cache similarities is beneficial for the model performance. This is no surprise, as the value cache directly contributes to the output vector of the attention layer. Although our experiments utilize $\mathit{L_{0}}$ masks to accelerate training process, model transformation can benefit any MHA to GQA conversion process.
Figure 4: The shared sparsity of $\mathit{L_{0}}$ masks across blocks allows different pruning speeds for different blocks, leading to a more stable training process.
In addition, during the experiment, we found that the model with introduction of new KV heads performs much better than the model retaining original ones. That’s why we choose not to retain any original KV heads, this setting also allows different pruning speeds for different blocks. Figure 4 shows the actual average size of masks in each block at different target sizes.
@@ 5 Conclusion
In this paper, we propose a general method for pruning an MHA model into a GQA model with any compression ratio of key-value heads. We find that applying appropriate orthogonal transformations to the model can increase the similarity between key-value heads without changing the model, thereby reducing the difficulty of model pruning. Furthermore, we introduce $\mathit{L_{0}}$ masks during pruning training, which reduce the impact of directly eliminating parameters on the model. Our method is applicable to all key-value head pruning conditions.
@@
Our work has two main limitations. First, we don’t delve into the grouping method, and current grouping method may not be the optimal one. How to find a more reasonable grouping method is one of the future research directions. Moreover, our method entirely relies on the statistical mathematical features of each attention head, without considering semantic information of each attention head. In fact, compressing attention heads based on semantic information is also a promising direction Tang et al. (2024).
@@ Appendix B Details of datasets
The sizes of sub datasets are shown in Table  3.
@@
Arc-C Arc-E OpenbookQA Which color shirt will reflect the most light on a hot, sunny day? Choices: [’black’, ’blue’, ’red’, ’white’] Answer:
HellaSwag Please choose the most appropriate text to complete the passage below:
@@
BoolQ The Coroner – The BBC announced on 2 March 2017 that there would be no further series.
@@
Winogrande Choose the most sensible text to replace the ’_’ in the following sentence: Natalie was less religous than Patricia, therefore _ attended church services more often on Sundays.
@@
PIQA Goal: how do you flood a room?
@@
SIQA Sasha took him to vegas for a vacation.

### v2

@@ 4.4 Analysis of the results
Although our experiments utilize $\mathit{L_{0}}$ regularization to accelerate the training process, model transformation can benefit any MHA to GQA conversion process. In certain cases, value-based grouping also contributed to result improvements. For experiments where grouping failed to yield performance gains, we hypothesize that cumulative errors introduced by the transformation may have a negative impact on model performance.
Figure 3: The shared sparsity of $\mathit{L_{0}}$ masks across blocks allows different pruning speeds for different blocks in LLaMA2-7B, leading to a more stable training process.
In addition, during the experiments, we found that the model retaining original KV heads failed to converge during the pruning process. That’s why we choose not to retain any original KV heads, this setting also allows different pruning speeds for different blocks. Figure 3 shows the actual average size of masks in each block at different target sizes. Attention heads closer to the input layer are pruned last, as pruning these heads significantly impacts all subsequent layers.
@@ 5 Conclusion
In this paper, we propose a general method for converting an MHA model into a GQA model with any compression ratio of KV heads. We find that applying orthogonal transformations to attention heads based on Procrustes analysis can enhance the similarity between KV heads without changing the model, thereby improving its performance after pruning. Furthermore, we introduce $\mathit{L_{0}}$ regularization during pruning training, which reduces the impact of directly eliminating parameters on the model. Our method is applicable to all KV head pruning conditions.
@@
Our work has two main limitations. First, we do not delve into the grouping method, and the current approach can be further optimized. Identifying a more effective grouping strategy is one of the potential directions for future research. Moreover, our method entirely relies on the statistical mathematical features of attention heads, without considering their semantic information. In fact, compressing attention heads based on semantic information is also a promising direction Tang et al. (2024).
@@ Appendix B Details of datasets
The sizes of sub datasets are shown in Table  6.
@@
Arc-C
Arc-E
OpenbookQA
Which color shirt will reflect the most light on a hot, sunny day? Choices: [’black’, ’blue’, ’red’, ’white’] Answer:
HellaSwag
Please choose the most appropriate text to complete the passage below:
@@
BoolQ
The Coroner – The BBC announced on 2 March 2017 that there would be no further series.
@@
Winogrande
Choose the most sensible text to replace the ’_’ in the following sentence: Natalie was less religous than Patricia, therefore _ attended church services more often on Sundays.
@@
PIQA
Goal: how do you flood a room?
@@
SIQA
Sasha took him to vegas for a vacation.
@@
Appendix C Time cost of model calibration and transformation
model LLaMA-2-7B Sheared-llama-1.3B
calibration 5min 3min
GQA16
transformation
w/ grouping 22min \
w/o grouping 4min \
GQA8
transformation
w/ grouping 37min 20min
w/o grouping 17min 2min
GQA4
transformation
w/ grouping 1h2min 35min
w/o grouping 29min 14min
Table 6: Time cost of model calibration and transformation. It requires at most one hour to complete this procedure. All calculations are performed by 1 A100 GPU.